OpenStudy (yanasidlinskiy):

B:)) "Determine the mean and standard deviation of the random variable X. Round your results to the nearest penny."

hartnn (hartnn):

do you know the formula?

OpenStudy (yanasidlinskiy):

Yes it's like... Standard Deviation over the sample...

OpenStudy (yanasidlinskiy):

Unless I'm wrong...

hartnn (hartnn):

its same what do you get for mean?

hartnn (hartnn):

\(\mu = mean = \sum p_ix_i \\ \sigma = \sqrt{\dfrac{1}{N}\sum (x_i-\mu )^2}\)

OpenStudy (yanasidlinskiy):

N would be what tho?...

hartnn (hartnn):

number of samples, here it would be = 2

OpenStudy (yanasidlinskiy):

Ahh ok cuz we had 2 samples or whatever right?

hartnn (hartnn):

yes. mean = 35*P(35) + (-1)P(-1) = ... ?

hartnn (hartnn):

use the values we got from part (a)

OpenStudy (yanasidlinskiy):

Yea, that's what I'm doing rn

OpenStudy (yanasidlinskiy):

Okay, I'm not sure if I'm doing this right but I'm just gonna plug in the numbers, cuz I'm not confident about it..... 35*.0263(35)+(-1).9737(-1)

hartnn (hartnn):

why 35 and -1 are multiplied twice?

OpenStudy (yanasidlinskiy):

Oh my bad lol

hartnn (hartnn):

ohh.. P(35) is probability of 35, not probability*35

hartnn (hartnn):

so, mean is 35*(0.0263) + (-1)(0.9737) = ...

OpenStudy (yanasidlinskiy):

So, 1.9?

hartnn (hartnn):

i get -0.0526...

OpenStudy (yanasidlinskiy):

What?

OpenStudy (yanasidlinskiy):

Desktop calculator... My laptop lol

hartnn (hartnn):

now start calculating std deviation

OpenStudy (yanasidlinskiy):

Okay, my sample would be the same, right?

hartnn (hartnn):

yes same

OpenStudy (yanasidlinskiy):

Same formula as well, right?

OpenStudy (yanasidlinskiy):

Like.... Wait. Actually how would I calculate it.

hartnn (hartnn):

This: \( \\ \sigma = \sqrt{\dfrac{1}{N}\sum (x_i-\mu )^2}\)

hartnn (hartnn):

N = 2, x1 = 35 x2 = -1 mu = -0.0526

hartnn (hartnn):

\(\sqrt {\dfrac{1}{2} [(35-(-0.0526))^2+(-1-(0.0526))^2]}\) just substituted the values

OpenStudy (yanasidlinskiy):

24.8?

OpenStudy (yanasidlinskiy):

Gee, Okay, I may need to remember this ahha

hartnn (hartnn):

the formulas, yes.

OpenStudy (yanasidlinskiy):

No, I know the formulas just messing with the numbers is really insane.

hartnn (hartnn):

good luck!

OpenStudy (yanasidlinskiy):

Ughgh, yes thank you